OpenStudy (anonymous):

simplify f(x)= (ln(2x-1)*(x+2)^3)\(1-4x)^2

OpenStudy (paxpolaris):

\[f(x)= {\ln \left( 2x-1 \right)\left( x+2 \right)^3 \over \left( 1-4x \right)^2}\] why can't we leave it as it is?